# A Hitchhiker’s Guide to Understanding Performances of Two-Class Classifiers

### v1

@@ 4 Conclusion
In this paper, we introduced a practical hitchhiker’s guide to understanding the performance of two-class classifiers. We organized our guide into four distinct scenarios, each corresponding to a specific user profile, covering the theoretical analyst, the method designer, the benchmarker, and the application developer. For each scenario, we provided practical examples of challenges each user may encounter when evaluating classifier performance and identified the most suitable tools for their needs. Central to our approach is the adaptation of the Tile, which arranges an infinite number of ranking scores into a 2D map. By leveraging different mappings, we demonstrated the versatility of the Tile and its various flavors to address the diverse requirements of these user profiles. This guide offers a flexible and robust framework that enables users to effectively evaluate, rank, and interpret the performance of two-class classifiers, making it a valuable resource for researchers, practitioners, and developers alike.
@@ Appendix A Supplementary Material
We provide in this supplementary material (1) the list of symbols used in this paper (Section A.1), (2) a description of the software in the Python Jupyter Notebook (Sec. A.2), (3) a definition of the Relative-Skill Tile (Sec. A.3), (4) more details about the illustration used in the paper (Sec. A.4), (5) an analysis of the behavior of scores, using the Correlation Tile, explaining the patterns observed in the various tiles presented in the paper (Sec. A.5), and finally (6) the comprehensive report generated by the Python Jupyter Notebook for our illustration (Sec. A.6).
@@ A.2 Software Description
The Python Jupyter Notebook generates a comprehensive report related to the illustration of this paper, featuring all the tiles presented in the paper and more. The various tiles are obtained by integrating two types of information. The first involves point-specific data, such as the value of a ranking score, a rank, an entity, or a correlation value, as detailed in this paper. This is achieved by discretizing the Tile using a grid size parameter along the two axis, $a$ and $b$, and mapping the respective data, accordingly to the ranking scores in the Tile, to a point $(a,b)$. The second type, relevant when priors are fixed (i.e., when all compared performances have the same class priors), pertains to area-based information. The algorithm used to identify these areas of interest, described in paper B [45], determines the areas (and their respective boundaries) within the Tile where an entity holds a specific rank (typically, the first or the last) without requiring a discretization of the Tile. The hatched areas on the various tiles provided in the paper (e.g., the Value Tile or Ranking Tile) are drawn using this algorithm, as well as the white boundary lines on the Baseline Value Tile and the State-of-the-Art Value Tile.
@@ A.4 More Details on the Illustration
A.5 Behavior of Scores
The Correlation Tile allows to depict the behavior of any score, showing the rank correlations between that score and all canonical ranking scores, for a given performance distribution. On Fig. 10, we analyze the rank correlation, using Spearman correlation coefficient (Spearman’s $\rho$), for $6$ scores belonging to the ranking scores (namely $TNR$, $TPR$, $NPV$, $PPV$, $A$, and $F_{1}$) and $3$ distributions of performances (all $3$ distributions are uniform but are on different sets of performances). We analyze these scores (1) for a uniform distribution over all performances, (2) for a uniform distribution of performances having fixed priors corresponding to those of BDD100K, and (3) for a uniform distribution of performances, where the set of performances is the one of our illustration. Comparing the figures on the left with those on the center, we see the effect of having fixed priors with $\pi_{+}=0.1242$. The figures on the right show the effect of having fixed priors and an arbitrary choice of performances (the ones of the $74$ state-of-the-art semantic segmentation models). These last two conditions seems to be responsible for the patterns (horizontal bands) observed in the various tiles of the paper.
@@ A.5 Behavior of Scores
A.6 Report Generated by the Python Jupyter Notebook for our Illustration
@@
ANN ♠ 0.8390 0.0368 0.0368 0.1064
ANN ♥ 0.8615 0.0142 0.0142 0.1106
ANN ♠ 0.8713 0.0045 0.0045 0.0987
APCNet ♠ 0.8365 0.0393 0.0393 0.1010
APCNet ♥ 0.8599 0.0159 0.0159 0.1114
BiSeNetV1 ♠ 0.8453 0.0304 0.0304 0.0992
BiSeNetV1 ♣ 0.8357 0.0401 0.0401 0.1159
BiSeNetV2 ♠ 0.8332 0.0426 0.0426 0.1088
CCNet ♠ 0.7896 0.0862 0.0862 0.1052
CCNet ♥ 0.8621 0.0136 0.0136 0.1101
CCNet ♠ 0.8719 0.0039 0.0039 0.0990
CGNet ♠ 0.6831 0.1926 0.1926 0.1172
DANet ♠ 0.8030 0.0728 0.0728 0.1069
DANet ♥ 0.8626 0.0131 0.0131 0.1107
DANet ♠ 0.8717 0.0040 0.0040 0.0978
DeepLabV3+ ♠ 0.8175 0.0583 0.0583 0.1042
DeepLabV3+ ♥ 0.8627 0.0131 0.0131 0.1106
DeepLabV3+ ♠ 0.8718 0.0040 0.0040 0.0994
DeepLabV3 ♠ 0.8424 0.0333 0.0333 0.1029
DeepLabV3 ♥ 0.8633 0.0125 0.0125 0.1102
DeepLabV3 ♠ 0.8710 0.0048 0.0048 0.1002
DeepLabV3 ♣ 0.8609 0.0149 0.0149 0.1164
DMNet ♠ 0.8517 0.0240 0.0240 0.0937
DMNet ♥ 0.8618 0.0140 0.0140 0.1107
DNLNet ♠ 0.8316 0.0441 0.0441 0.1038
DNLNet ♥ 0.8633 0.0124 0.0124 0.1099
DPT ♥ 0.8635 0.0123 0.0123 0.1092
EMANet ♠ 0.8383 0.0375 0.0375 0.0996
EncNet ♠ 0.8437 0.0320 0.0320 0.0960
EncNet ♥ 0.8633 0.0125 0.0125 0.1093
ERFNet ♠ 0.8450 0.0308 0.0308 0.1073
Fast-SCNN ♠ 0.8290 0.0468 0.0468 0.1106
FastFCN ♠ 0.8397 0.0361 0.0361 0.0989
@@
FCN ♠ 0.8476 0.0282 0.0282 0.0977
FCN ♥ 0.8657 0.0100 0.0100 0.1079
FCN ♠ 0.8697 0.0061 0.0061 0.1022
GCNet ♠ 0.8104 0.0654 0.0654 0.1047
GCNet ♥ 0.8635 0.0123 0.0123 0.1090
GCNet ♠ 0.8715 0.0043 0.0043 0.0989
ICNet ♠ 0.8586 0.0172 0.0172 0.1135
ISANet ♠ 0.8304 0.0454 0.0454 0.1038
ISANet ♥ 0.8569 0.0189 0.0189 0.1123
ISANet ♠ 0.8720 0.0038 0.0038 0.0987
K-Net ♥ 0.8630 0.0127 0.0127 0.1144
Mask2Former ♠ 0.8621 0.0137 0.0137 0.1186
Mask2Former ♥ 0.8661 0.0097 0.0097 0.1149
MaskFormer ♥ 0.8624 0.0133 0.0133 0.1139
NonLocal Net ♠ 0.8269 0.0489 0.0489 0.1067
NonLocal Net ♥ 0.8649 0.0109 0.0109 0.1092
NonLocal Net ♠ 0.8716 0.0042 0.0042 0.0992
OCRNet ♠ 0.8523 0.0235 0.0235 0.1132
OCRNet ♥ 0.8635 0.0123 0.0123 0.1112
OCRNet ♠ 0.8707 0.0051 0.0051 0.0999
PointRend ♠ 0.8422 0.0335 0.0335 0.0914
PointRend ♥ 0.8611 0.0147 0.0147 0.1101
PSANet ♠ 0.8268 0.0489 0.0489 0.1088
PSANet ♥ 0.8634 0.0123 0.0123 0.1087
PSPNet ♠ 0.8528 0.0230 0.0230 0.0970
PSPNet ♥ 0.8629 0.0129 0.0129 0.1100
PSPNet ♠ 0.8710 0.0048 0.0048 0.1011
PSPNet ♣ 0.8584 0.0174 0.0174 0.1165
SAN ♣ 0.8544 0.0214 0.0214 0.1155
SegFormer ♠ 0.8670 0.0088 0.0088 0.1135
SegFormer ♥ 0.8637 0.0121 0.0121 0.1115
Segmenter ♥ 0.8604 0.0154 0.0154 0.1150
Semantic FPN ♠ 0.8600 0.0158 0.0158 0.0838
Semantic FPN ♥ 0.8225 0.0533 0.0533 0.0851
SETR ♠ 0.8663 0.0095 0.0095 0.1168
SETR ♥ 0.8602 0.0155 0.0155 0.1128
STDC ♠ 0.8411 0.0347 0.0347 0.1078
UPerNet ♠ 0.8344 0.0413 0.0413 0.1096
UPerNet ♥ 0.8619 0.0138 0.0138 0.1103
UPerNet ♠ 0.8718 0.0039 0.0039 0.0983
[Uncaptioned image][Uncaptioned image]
@@
  • SegFormer ♠ (max rank: 14, mean rank: 6.8558)
  • SETR ♠ (max rank: 14, mean rank: 4.0495)

### v3

@@ 5 Conclusion
In this paper, we introduced a hitchhiker’s guide to understanding the performance of two-class classifiers. We organized our guide into four distinct scenarios, each corresponding to a specific user profile, covering the theoretical analyst [Uncaptioned image], the method designer [Uncaptioned image], the benchmarker [Uncaptioned image], and the application developer [Uncaptioned image]. For each scenario, we provided examples of challenges each user may encounter when comparing classifier performances and identified the most suitable tools for their needs. Central to our approach is the idea of displaying different types of information, called flavors, on the Tile. By leveraging different mappings, we demonstrated the versatility of the Tile and its various flavors to address the diverse requirements of these user profiles. This guide offers a flexible and robust framework that enables users to effectively evaluate, rank, and interpret the performance of two-class classifiers, making it a valuable resource for researchers, practitioners, and developers alike.
@@ Appendix A Supplementary Material
We provide in this supplementary material (1) the list of symbols used in this paper (Section A.1), (2) a description of the software in the Python Jupyter Notebook (Sec. A.2), (3) a definition of the Relative-Skill Tile (Sec. A.3), (4) more details about the illustration used in the paper (Sec. A.4), (5) an analysis of the influence of a learning set (Sec. A.5), (6) an analysis of the behavior of scores, using the Correlation Tile, explaining the patterns observed in the various tiles presented in the paper (Sec. A.6), and finally (7) the comprehensive report generated by the Python Jupyter Notebook for our illustration (Sec. A.7).
@@ A.2 Software Description
The Python Jupyter Notebook generates a comprehensive report related to the illustration of this paper, featuring all the tiles presented in the paper and more. It takes as input a list of two-class classification performances (i.e., the values of $P(\{\omega\})$, for all $\omega\in\{tn,fp,fn,tp\}$). The various tiles are obtained by integrating two types of information. The first involves point-specific data, such as the value of a ranking score, a rank, an entity, or a correlation value, as detailed in this paper. This is achieved by discretizing the Tile using a grid size parameter along the two axis, $a$ and $b$, and mapping the respective data, accordingly to the ranking scores in the Tile, to a point $(a,b)$. The second type, relevant when priors are fixed (i.e., when all compared performances have the same class priors), pertains to area-based information. The algorithm used to identify these areas of interest, described in [36], determines the areas (and their respective boundaries) within the Tile where an entity holds a specific rank (typically, the first or the last) without requiring a discretization of the Tile. The hatched areas on the various tiles provided in the paper (e.g., the Value Tile or Ranking Tile) are drawn using this algorithm, as well as the white boundary lines on the Baseline Value Tile and the State-of-the-Art Value Tile.
A.2.1 Three Methods to Compute Value Tiles
In practice, several methods can be used to compute the Value Tile, depending on the availability of data.
  1. 1.
    Direct computation: If the values of $P(\{fp\})$, $P(\{fn\})$, $P(\{tp\})$, and $P(\{tn\})$ are known, the first method consists in computing the values of the ranking scores in each point of the Tile using Eq. 3.
  2. 2.
    Interpolation: Knowing the values taken by the scores $TPR$, $TNR$, $NPV$, and $PPV$ at the four corners of the Tile, the values taken by the remaining scores can be determined by averaging, using $f$-means, vertically with $f:x\mapsto x^{-1}$ (i.e. the harmonic mean) and horizontally with $f:x\mapsto(1-x)^{-1}$ [36]. As one can see, the non-linearities differ between the vertical and horizontal axes.
  3. 3.
    Equation system resolution: The third method involves first, determining the values of $P(\{fp\})$, $P(\{fn\})$, $P(\{tp\})$, and $P(\{tn\})$ by solving a system of $4$ equations with $4$ unknowns, then, computing the values of the ranking scores using Eq. 3. Specifically, if the values of three ranking scores from the Value Tile are known, or if two ranking scores and the priors (for fixed priors) with $\pi_{-}=P(\{tn\})+P(\{fp\})$ and $\pi_{+}=P(\{fn\})+P(\{tp\})$, then, knowing that $P(\{fp\})+P(\{fn\})+P(\{tp\})+P(\{tn\})=1$ makes the system complete.
These three methods require discretizing the Tile. This can be implemented with a grid size parameter, which defines linearly spaced values for the two axes, $a$ and $b$. The provided code implements the direct computation method.
@@ A.4 More Details on the Illustration
A.5 Influence of a Learning Set
The influence of a learning set on the ranking of a semantic segmentation model can be seen by placing the Ranking Tiles corresponding to that model trained on the different learning sets side by side. As shown in Fig. 10, training PSPNet on ♠ Pascal VOC 2012 and ♣ COCO-Stuff 164k performs better than training it on ♠ Cityscapes. But the best stability over the tile is achieved by training PSPNet on ♥ ADE20K.
Figure 10: Ranking Tiles showing the rank of PSPNet trained on different learning datasets (from left to right: ♠ Cityscapes, ♥ ADE20K, ♠ Pascal VOC 2012, and ♣ COCO-Stuff 164k). We observe that the learning dataset strongly influences the ranking behavior.
A.6 Behavior of Scores
The Correlation Tile allows to depict the behavior of any score, showing the rank correlations between that score and all canonical ranking scores, for a given performance distribution. On Fig. 11, we analyze the rank correlation, using Spearman correlation coefficient (Spearman’s $\rho$), for $6$ scores belonging to the ranking scores (namely $TNR$, $TPR$, $NPV$, $PPV$, $A$, and $F_{1}$) and $3$ distributions of performances (all $3$ distributions are uniform but are on different sets of performances). We analyze these scores (1) for a uniform distribution over all performances, (2) for a uniform distribution of performances having fixed priors corresponding to those of BDD100K, and (3) for a uniform distribution of performances, where the set of performances is the one of our illustration. Comparing the figures on the left with those on the center, we see the effect of having fixed priors with $\pi_{+}=0.1242$. The figures on the right show the effect of having fixed priors and an arbitrary choice of performances (the ones of the $74$ state-of-the-art semantic segmentation models). For the explanation of the quasi-horizontal bands, see Tab. 2.
@@ A.6 Behavior of Scores
Set 1 Set 2 Set 3
Num. perfs. (size) $1000$ $1000$ $74$
$\min(\theta)$ $0.000628$ $0.0000634$ $0.0971$
$\max(\theta)$ $1.57$ $1.56$ $0.170$
$\mu(\theta)$ $0.806$ $0.242$ $0.125$
$\sigma(\theta)$ $0.496$ $0.281$ $0.00997$
Table 2: Comparison of the three sets of two-class classification performances used in Fig. 11: (1) a uniform distribution over all possible performances, (2) over all performances with a prior of the positive class equal to the one in our illustration ($\pi_{+}=0.124227$), and (3) over the $74$ performances compared in our illustration. The score $\theta$ is given by $\theta=\arctan(\frac{P(\{tp\})}{P(\{tn\})})$. The set for which the standard deviation $\sigma(\theta)$ is minimal is the one for which $P(\{tn\})$ is the closer to be proportional to $P(\{tp\})$. It is the third one. This is why one can observe quasi horizontal bands in the Correlation Tiles on the right hand side of Fig. 11.
A.7 Report Generated by the Python Jupyter Notebook for our Illustration
@@
ANN ♠ 0.8390 0.0368 0.0178 0.1064
ANN ♥ 0.8615 0.0142 0.0136 0.1106
ANN ♠ 0.8713 0.0045 0.0255 0.0987
APCNet ♠ 0.8365 0.0393 0.0232 0.1010
APCNet ♥ 0.8599 0.0159 0.0128 0.1114
BiSeNetV1 ♠ 0.8453 0.0304 0.0250 0.0992
BiSeNetV1 ♣ 0.8357 0.0401 0.0083 0.1159
BiSeNetV2 ♠ 0.8332 0.0426 0.0154 0.1088
CCNet ♠ 0.7896 0.0862 0.0190 0.1052
CCNet ♥ 0.8621 0.0136 0.0142 0.1101
CCNet ♠ 0.8719 0.0039 0.0253 0.0990
CGNet ♠ 0.6831 0.1926 0.0070 0.1172
DANet ♠ 0.8030 0.0728 0.0173 0.1069
DANet ♥ 0.8626 0.0131 0.0136 0.1107
DANet ♠ 0.8717 0.0040 0.0265 0.0978
DeepLabV3+ ♠ 0.8175 0.0583 0.0200 0.1042
DeepLabV3+ ♥ 0.8627 0.0131 0.0136 0.1106
DeepLabV3+ ♠ 0.8718 0.0040 0.0248 0.0994
DeepLabV3 ♠ 0.8424 0.0333 0.0213 0.1029
DeepLabV3 ♥ 0.8633 0.0125 0.0140 0.1102
DeepLabV3 ♠ 0.8710 0.0048 0.0241 0.1002
DeepLabV3 ♣ 0.8609 0.0149 0.0078 0.1164
DMNet ♠ 0.8517 0.0240 0.0305 0.0937
DMNet ♥ 0.8618 0.0140 0.0135 0.1107
DNLNet ♠ 0.8316 0.0441 0.0205 0.1038
DNLNet ♥ 0.8633 0.0124 0.0143 0.1099
DPT ♥ 0.8635 0.0123 0.0150 0.1092
EMANet ♠ 0.8383 0.0375 0.0246 0.0996
EncNet ♠ 0.8437 0.0320 0.0282 0.0960
EncNet ♥ 0.8633 0.0125 0.0149 0.1093
ERFNet ♠ 0.8450 0.0308 0.0169 0.1073
Fast-SCNN ♠ 0.8290 0.0468 0.0136 0.1106
FastFCN ♠ 0.8397 0.0361 0.0253 0.0989
@@
FCN ♠ 0.8476 0.0282 0.0265 0.0977
FCN ♥ 0.8657 0.0100 0.0163 0.1079
FCN ♠ 0.8697 0.0061 0.0221 0.1022
GCNet ♠ 0.8104 0.0654 0.0196 0.1047
GCNet ♥ 0.8635 0.0123 0.0152 0.1090
GCNet ♠ 0.8715 0.0043 0.0253 0.0989
ICNet ♠ 0.8586 0.0172 0.0107 0.1135
ISANet ♠ 0.8304 0.0454 0.0204 0.1038
ISANet ♥ 0.8569 0.0189 0.0120 0.1123
ISANet ♠ 0.8720 0.0038 0.0256 0.0987
K-Net ♥ 0.8630 0.0127 0.0098 0.1144
Mask2Former ♠ 0.8621 0.0137 0.0056 0.1186
Mask2Former ♥ 0.8661 0.0097 0.0093 0.1149
MaskFormer ♥ 0.8624 0.0133 0.0103 0.1139
NonLocal Net ♠ 0.8269 0.0489 0.0176 0.1067
NonLocal Net ♥ 0.8649 0.0109 0.0150 0.1092
NonLocal Net ♠ 0.8716 0.0042 0.0250 0.0992
OCRNet ♠ 0.8523 0.0235 0.0110 0.1132
OCRNet ♥ 0.8635 0.0123 0.0130 0.1112
OCRNet ♠ 0.8707 0.0051 0.0243 0.0999
PointRend ♠ 0.8422 0.0335 0.0328 0.0914
PointRend ♥ 0.8611 0.0147 0.0142 0.1101
PSANet ♠ 0.8268 0.0489 0.0155 0.1088
PSANet ♥ 0.8634 0.0123 0.0155 0.1087
PSPNet ♠ 0.8528 0.0230 0.0273 0.0970
PSPNet ♥ 0.8629 0.0129 0.0142 0.1100
PSPNet ♠ 0.8710 0.0048 0.0231 0.1011
PSPNet ♣ 0.8584 0.0174 0.0078 0.1165
SAN ♣ 0.8544 0.0214 0.0088 0.1155
SegFormer ♠ 0.8670 0.0088 0.0107 0.1135
SegFormer ♥ 0.8637 0.0121 0.0127 0.1115
Segmenter ♥ 0.8604 0.0154 0.0092 0.1150
Semantic FPN ♠ 0.8600 0.0158 0.0404 0.0838
Semantic FPN ♥ 0.8225 0.0533 0.0391 0.0851
SETR ♠ 0.8663 0.0095 0.0074 0.1168
SETR ♥ 0.8602 0.0155 0.0115 0.1128
STDC ♠ 0.8411 0.0347 0.0164 0.1078
UPerNet ♠ 0.8344 0.0413 0.0146 0.1096
UPerNet ♥ 0.8619 0.0138 0.0139 0.1103
UPerNet ♠ 0.8718 0.0039 0.0259 0.0983
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]
@@
  • SegFormer ♠ (max rank: 14, mean rank: 6.85568)
  • SETR ♠ (max rank: 14, mean rank: 4.04938)